# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

### v1

@@ Abstract
Open-Vocabulary object detectors inherently generalize to an unrestricted set of categories, enabling recognition through simple textual prompting. However, adapting these models to rare classes or reinforcing their abilities on specialized domains remains essential. While recent methods rely on monolithic adaptation strategies with a single set of weights, we embrace modular deep learning. We introduce DitHub, a framework designed to build and maintain a library of efficient adaptation modules. Inspired by Version Control Systems, DitHub manages expert modules as branches that can be fetched and merged as needed. This modular approach allows us to conduct an in-depth exploration of the compositional properties of adaptation modules, marking the first such study in Object Detection. Our method achieves state-of-the-art performance on the ODinW-13 benchmark and ODinW-O, a newly introduced benchmark designed to assess class reappearance. For more details, visit our project page: https://aimagelab.github.io/DitHub/
@@ 1 Introduction
Figure 1: DitHub: a novel system to share and update deep adaptation modules for Open-Vocabulary Object Detection. [^2] The creature in the red box is an axolotl, symbolizing rarity as a deviation from pre-training. We use a stylized axolotl as the icon for DitHub.
Recent approaches, such as those detailed in [2], have extended vision-language detectors to incrementally accommodate new categories while preserving robust zero-shot capabilities [34]. However, these methods often employ monolithic adaptation, where all newly acquired knowledge is condensed into a single set of weights. This approach presents challenges in real-world scenarios that demand updates to specific concepts which may reappear under varied input modalities. For instance, in safety and security applications, the class “person” may need to be detected not only in standard RGB imagery but also in thermal imagery, necessitating continuous model adaptation to ensure consistent performance across different data types. Additionally, for rare or complex categories, incremental adaptation is crucial to refine the base model for fine-grained concepts. In such scenarios, a monolithic deep architecture faces challenges similar to those of managing a complex program written on a single page of code. Namely, when treating multiple classes within a unique set of weights, it becomes hard to update specific concepts selectively, without compromising information related to other concepts. Considering less-represented categories, the corresponding concepts can become blurred within the unique set of weights, leading to potential degradation in performance over time.
@@
With modularity at its core, we introduce DitHub (where D stands for Detection), a framework inspired by Version Control Systems that efficiently manages a growing library of detection modules. This library expands incrementally as new groups of categories, referred to as tasks [41], are introduced. Upon encountering a new task, DitHub initializes an efficient module (e.g., LoRA [11]), which first undergoes a broad pre-tuning phase across all classes of the task, referred to as the warmup phase. This stage establishes a solid and class-agnostic foundation, enabling the subsequent specialization of the module for specific classes. Following this phase, the module branches into specialized experts, each optimized independently. To enable selective updates for classes previously introduced to the model, DitHub first performs a fetch operation to access potentially existing class-specific modules. If these exist, instead of starting the training from the warmup, DitHub uses a merged initialization — combining the retrieved pre-existing class-specific modules with the current warmup module. This merged initialization facilitates smooth domain knowledge transfer and ensures continuity of class-specific learning across tasks.
Moreover, to address potential memory constraints, DitHub employs a subset of shared learnable parameters across different adaptation modules, enhancing efficiency without compromising individual class adaptability. At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and open-vocabulary capabilities. A high-level overview of our approach is depicted in Footnote 2.
In the context of [2], we evaluate DitHub on the ODinW-13 [17] benchmark, achieving state-of-the-art performance within the Incremental Vision-Language Object Detection paradigm. To further assess our proposal, we introduce ODinW-O (Overlapped), a curated subset of ODinW-35 [16] that focus on classes re-appearing across multiple tasks. Our simple yet effective methodology — without additional loss terms or architectural modifications — allows us to investigate the compositional capabilities of efficient modules, marking the first exploration of this kind in Object Detection. In summary, our key contributions are:
@@ 2 Related Works
Incremental Vision-Language Object Detection
Vision-Language Object Detection (VLOD), also referred to as Open-Vocabulary detection [2], is the task of detecting and recognizing objects beyond a fixed set of predefined categories. Early approaches leverage powerful visual and textual encoders (e.g., CLIP [34]) to enable object detectors to recognize virtually any object specified by a textual query [8, 49]. A fundamental trend in this field has been the reliance on large-scale pre-training on extensive datasets [37], which has become standard practice for tasks such as classification. In the context of Vision-Language Object Detection, state-of-the-art methods have followed this paradigm, with GLIP [17] being one of the first notable examples. GLIP reframes object detection as a phrase-grounding problem, leading to a pre-trained model capable of generalizing to unseen objects by integrating textual and visual semantics from large-scale datasets. Building on GLIP’s problem formulation, Grounding DINO [21] extends this paradigm by incorporating merged visual and textual semantics at multiple network stages.
Recently, Incremental Vision-Language Object Detection (IVLOD) [2] has emerged as a natural extension of the VLOD paradigm. By incorporating Incremental Learning [41, 14, 36, 45], IVLOD addresses the challenge of fine-tuning a pre-trained Vision-Language model on specific object categories while preserving its zero-shot [29] capabilities. Following the approach of [2], we adopt Grounding DINO as our backbone and utilize IVLOD as the foundation to explore the merging of class-specific modules.
Parameter Efficient Fine-Tuning and Model Merging
The rise of Transformer models [42, 5] has made pre-training a standard approach in machine learning. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix Tuning [18], Adapters [10], and LoRA [11], adapt pre-trained models to downstream tasks by updating only a small fraction of parameters while keeping the rest frozen.
In the context of Incremental Vision-Language Object Detection, ZiRa [2] introduces a distinct approach that can be viewed as a specialized form of parallel adapters [23], incorporating reparameterizable side branches. In contrast, our work adopts LoRA and explores the potential of merging class-specific LoRA modules.
Several studies have investigated LoRA module fusion. An early approach relies on simple arithmetic operations such as addition and subtraction [50]. More advanced techniques focus on learning optimal combination coefficients, determining the contribution of each LoRA module in the merging process [12, 47, 46, 48]. Recently, researchers have introduced closed-form solutions for LoRA merging [38] and proposed regularization objectives to ensure that the fused modules remain close to the pre-trained weight distribution [31]. In this respect, our approach takes a novel direction, being the first to study module merging in the context of object detection.
@@ 3 Preliminaries
Problem setting. We consider a sequence of tasks $\{1,\dots,T\}$, where each task $t$ corresponds to a dataset $D_{t}$. Each dataset consists of samples $(x,y)$, where $x$ represents the input data, and $y$ contains the associated labels, including ground truth bounding boxes.
@@
$\min_{\Delta}\mathbb{E}_{(x,y)\sim D_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
@@
LoRA. Low-Rank Adaptation [11] is a technique to significantly reduce the number of trainable parameters. Given a generic layer with weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ — with $r\ll\min(d,k)$ — such that the adapted weight $W^{\prime}$ is:
$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=BA.$ (2)
The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to fine-tuning the entire model.
@@ 4 DitHub
Figure 2: Overview of DitHub. Through three primitives — i.e., branch, fetch, merge — DitHub allows to update a library of class-specific modules for Open-Vocabulary Object Detection, ensuring a balanced integration of both past and present knowledge.
Version Control Systems are ubiquitous tools in computer science, renowned for their capacity to efficiently manage code repositories. Inspired by this capability, we introduce DitHub— a framework that operates akin to a version control system, facilitating the creation and continuous expansion of a library of efficient adaptation modules for vision-language object detectors. To develop and maintain class-specific expert modules, DitHub employs a straightforward yet effective strategy that requires no architectural modifications or additional loss terms, as detailed in Sec. 4.1.
In line with the authors of [2], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection. We employ LoRA as a fine-tuning strategy, proposing a novel and efficient pipeline that handles its low-rank matrices with two specific objectives. In our approach, matrix $A$ is dedicated to class-specificity, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency, as further discussed in Sec. 4.2. Finally, when a specialized task arises or more fine-grained adaptation is required (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec. 4.3. The pseudocode of DitHub is provided in Alg. 1.
@@ 4.1 Training Class-Specific Modular Detectors
To manage a library of class-specific modules, the approach is to train a tailored $A$ matrix for each class, allowing it to specialize accordingly. On the contrary, the matrix $B$ is shared across classes and continually fine-tuned while tasks progress. Thus, at task $t$, we train a total of $|\mathcal{C}_{t}|+1$ matrices for each layer that requires adaptation. The total includes a matrix for each of the $|\mathcal{C}_{t}|$ classes introduced in the current task, along with a single matrix for $B$.
However, in object detection, creating class-specific experts is non-trivial, as a single image may contain multiple bounding boxes spanning different categories, making it difficult to focus on a certain class. To tackle this, we adopt a stochastic training strategy for specializing the $A$ matrices: given an image with multiple objects from different classes, we randomly select one of the present classes and execute a training step on its corresponding class-expert module. However, in our preliminary experiments, this strategy proved ineffective for rare categories, which may appear only a few times in real-world scenarios. Relying solely on random selection prevents the $A$ matrix of rare classes from accessing enough training data.
@@
Warmup. At the onset of the new task, a single and shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig. 2, yellow arrow) to capture class-agnostic yet task-specific knowledge. This phase, spanning a predefined number of epochs, establishes a common initialization for the $A$ modules, which are set to undergo specialization in the subsequent stage. By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [27, 33]. Often associated with improved linear mode connectivity between models [7], these properties are leveraged by our approach to select and combine class-specific modules based on input textual queries, as detailed in Sec. 4.3. Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
Specialization. Once the warmup phase is finished, we begin with the specialization phase, the cornerstone of DitHub. To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig. 2), each dedicated to a single class. These branches undergo independent optimization, represented by a blue arrow in Fig. 2, to capture class-specific capabilities. In doing so, for each training image, we employ the stochastic strategy outlined at the beginning of Sec. 4.1 to randomly select a single expert for training.
Notably, the modular design of our approach effectively handles potential class overlaps across consecutive tasks. In real-world scenarios, classes may reappear in new domains with altered distributions or semantics. For instance, consider the class “dog”, originally learned from real-world camera images, later appearing in a task involving thermal imagery. The goal is not only to retain prior knowledge but also to develop a “dog” expert capable of recognizing the entity in both domains, accommodating any semantic shifts. In this case, the modularity of DitHub enables the selective adaptation of the “dog” module, while minimizing interference with other classes, such as the “axolotl” module.
To favor selective updates, DitHub performs a fetch operation (Fig. 2) to retrieve any previously stored version of the specialized module, such as $A_{\text{dog}}^{\text{old}}$. If found, this module is merged (merge in Fig. 2) with the current task’s warmup matrix $A_{\text{wu}}$. For instance, considering the class “dog”, the merge operation follows a weighted strategy:
Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase. Details on the impact of $\lambda_{A}$ are in Sec. 5.3. Crucially, our fetching process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information.
@@
Input: $T$ disjoint tasks each containing $|\mathcal{C}_{t}|$ classes; library $\mathcal{A}\leftarrow\{\}$ of class experts; global $B$.
@@
              $A_{c}^{\text{cur}}\leftarrow$ use Equation 3 $\triangleright$  Merge
@@
     $B\leftarrow$ use Equation 4
@@ 4.2 Bridging Modularity with Memory Efficiency
Figure 3: Memory requirements as the number of classes grows.
To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class. As shown in Fig. 3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows. Sharing the $B$ matrix reduces this requirement by half  [^3] In Sec. A.2, we present DitHub’s results when $B$ is not shared.. Indeed, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder model alone (line [Uncaptioned image]). This makes DitHub’s memory demands manageable for practical applications. Also, the memory efficiency is further optimized by fine-tuning only a limited number of layers (see App. A).
Due to the shared design, the matrix $B$ is more susceptible to interference and catastrophic forgetting [25]. Consequently, we implement a tailored mechanism that ensures smooth updates as tasks progress. Specifically, at the end of each task $t$, we update the matrix $B$ by merging the previous checkpoint, $B_{t-1}$, with the weights learned after the current task, $B^{\text{opt}}$. This process is formulated as follows:
@@
Unlike Eq. 3, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state. This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, which will be further explained in Sec. 5.3.
4.3 Detection with Class-Specific Modules
With our library of efficient, class-specific modules in place for adapting the Vision-Language backbone, two disjoint inference routes are available for users. Given a textual query specifying the classes of interest, our model tokenizes the prompt and searches token by token to determine if corresponding expert modules are present in the library. If there are correspondences, those relevant modules are fused using a simple average to create a single composite adaptation module. Conversely, for requests targeting individual classes where fine-grained adaptation is desired, the corresponding module can be directly selected and deployed. In our experimental studies, we present results and discussions that address both scenarios.

### v2

@@ Abstract
Open-Vocabulary object detectors can generalize to an unrestricted set of categories through simple textual prompting. However, adapting these models to rare classes or reinforcing their abilities on multiple specialized domains remains essential. While recent methods rely on monolithic adaptation strategies with a single set of weights, we embrace modular deep learning. We introduce DitHub, a framework designed to build and maintain a library of efficient adaptation modules. Inspired by Version Control Systems, DitHub manages expert modules as branches that can be fetched and merged as needed. This modular approach allows us to conduct an in-depth exploration of the compositional properties of adaptation modules, marking the first such study in Object Detection. Our method achieves state-of-the-art performance on the ODinW-13 benchmark and ODinW-O, a newly introduced benchmark designed to assess class reappearance.
@@ 1 Introduction
Recent approaches, such as those detailed in [4], have extended Vision-Language detectors to incrementally accommodate new categories while preserving robust capabilities [39]. However, these methods employ monolithic adaptation, where all newly acquired knowledge is condensed into a single set of weights. This approach presents challenges in real-world scenarios that demand updates to specific concepts which may reappear under varied input modalities. For instance, in safety and security applications, the class “person” may need to be detected not only in standard RGB imagery but also in thermal imagery, necessitating continuous model adaptation to ensure consistent performance across different data types. Additionally, for rare or complex categories, incremental adaptation is crucial to refine the base model for fine-grained concepts. In such scenarios, a monolithic deep architecture faces challenges similar to managing a complex program written on a single page of code. Namely, when treating multiple classes within a unique set of weights, it becomes hard to update specific concepts selectively, without compromising information related to other concepts. Considering less-represented categories, the corresponding concepts can become blurred within the unique set of weights, leading to potential degradation in performance over time.
Figure 1: DitHub: a novel system to share and update efficient adaptation modules for Open-Vocabulary Object Detection. [^2] The creature in the red box is an axolotl, symbolizing rarity as a deviation from pre-training. We use a stylized axolotl as the icon for DitHub.
@@
With modularity at its core, we introduce DitHub (where D stands for Detection), a framework inspired by Version Control Systems that efficiently manages a growing library of detection modules. This library expands incrementally as new groups of categories, referred to as tasks [46], are introduced. Upon encountering a new task, DitHub initializes an efficient module (e.g., LoRA [13]), which first undergoes a broad pre-tuning phase across all classes of the task, referred to as the warmup phase. This stage establishes a solid and class-agnostic foundation, enabling the subsequent specialization of the module for specific classes. Following the warmup phase, the module branches into specialized experts, each optimized independently. To enable selective updates for classes previously introduced to the model, DitHub first performs a fetch operation to access potentially existing class-specific modules. If these exist, instead of starting the training from the warmup, DitHub uses a merged initialization — combining the retrieved pre-existing class-specific modules with the current warmup module. This merged initialization facilitates smooth domain knowledge transfer and ensures continuity of class-specific learning across tasks.
Moreover, to ensure scalability with respect to the number of classes, DitHub shares a subset of parameters across different adaptation modules, enhancing memory efficiency without compromising individual class adaptability. At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and Open-Vocabulary capabilities. A high-level overview of our approach is depicted in Footnote 2.
Following [4], we evaluate DitHub on the ODinW-13 [20] benchmark, where it outperforms the state-of-the-art by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP in the Incremental Vision-Language Object Detection setting. We also introduce ODinW-O (Overlapped), a curated subset of ODinW-35 [19] focusing on classes re-appearing across multiple tasks. Furthermore, we thoroughly investigate and ablate the compositional capabilities of our modular approach, marking the first such exploration in Object Detection. In summary, our key contributions are:
  • We investigate the importance of decoupling a warmup phase from a specialization one in the context of efficient adaptation modules.
@@ 2 Related Works
Incremental Vision-Language Object Detection. Vision-Language Object Detection (VLOD), also referred to as Open-Vocabulary detection [4], is the task of detecting and recognizing objects beyond a fixed set of predefined categories. Early approaches leverage powerful visual and textual encoders (e.g., CLIP [39]) to enable object detectors to recognize virtually any object specified by a textual query [10, 54]. A fundamental trend in this field has been the reliance on large-scale pre-training on extensive datasets [42], which has become standard practice for tasks such as classification. In the context of Vision-Language Object Detection, state-of-the-art methods have followed this paradigm, with GLIP [20] being one of the first notable examples. GLIP reframes Object Detection as a phrase-grounding problem, leading to a pre-trained model capable of generalizing to unseen objects by integrating textual and visual semantics from large-scale datasets. Building on GLIP’s problem formulation, Grounding DINO [24] extends this paradigm by incorporating merged visual and textual semantics at multiple network stages.
Recently, Incremental Vision-Language Object Detection (IVLOD) [4] has emerged as a natural extension of the VLOD paradigm. By incorporating Incremental Learning [46, 17, 41, 50], IVLOD addresses the challenge of fine-tuning a pre-trained Vision-Language model on specific object categories while preserving its zero-shot [34] capabilities. Following [4], we adopt Grounding DINO as the backbone and use IVLOD to explore the merging of class-specific modules.
Parameter Efficient Fine-Tuning and Model Merging. The rise of Transformer models [47, 7] has made pre-training a standard approach in machine learning. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix Tuning [21], Adapters [12], and LoRA [13], adapt pre-trained models to downstream tasks by updating only a small fraction of parameters while keeping the rest frozen.
In the context of IVLOD, ZiRa [4] introduces a distinct approach that can be viewed as a specialized form of parallel adapters [27], incorporating reparameterizable side branches. In contrast, our work adopts LoRA and explores the potential of merging class-specific LoRA modules.
Several studies have investigated LoRA module fusion. An early approach relies on simple arithmetic operations such as addition and subtraction [55]. More advanced techniques focus on learning optimal combination coefficients, determining the contribution of each LoRA module in the merging process [14, 52, 51, 28, 53]. Recently, researchers have introduced closed-form solutions for LoRA merging [43] and proposed regularization objectives to ensure that the fused modules remain close to the pre-trained weight distribution [36]. In this respect, our approach takes a novel direction, being the first to study module merging in the context of Object Detection.
@@ 3 Preliminaries
Problem setting. We consider a sequence of tasks $\{1,\dots,T\}$, where each task $t$ corresponds to a dataset $\mathcal{D}_{t}$. Each dataset consists of samples $(x,y)$, where $x$ represents the input data, and $y$ contains the associated labels, including ground truth bounding boxes.
@@
$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
@@
LoRA. Given a layer with weight $W\in\mathbb{R}^{d\times k}$, Low-Rank Adaptation [13] (LoRA) uses low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ – with $r\ll\min(d,k)$ – such that the adapted weight $W^{\prime}=W+\Delta W=W+BA$. The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to full fine-tuning.
@@ 4 DitHub
Figure 2: Overview of DitHub. Through three primitives – i.e., branch, fetch, merge – DitHub allows to update a library of class-specific modules for Open-Vocabulary Object Detection.
Version Control Systems are ubiquitous tools used to manage and track changes in code repositories. Inspired by this, we propose DitHub — a framework that operates akin to a Version Control System, enabling the creation and expansion of a library of efficient adaptation modules for Vision-Language object detectors. DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Sec. 4.1).
Following [4], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection and use LoRA for fine-tuning. We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Sec. 4.2). Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec. 4.3. The pseudocode of DitHub is provided in Alg. 1.
@@ 4.1 Training Class-Specific Modular Detectors
To manage a library of class-specific modules, our approach is to train a tailored $A$ matrix for each class to enable specialization. In contrast, the $B$ matrix is shared across classes and continuously fine-tuned as tasks progress. Thus, at task $t$, we train a total of $|\mathcal{C}_{t}|+1$ matrices per adaptable layer — one for each class and one shared $B$ matrix.
However, in Object Detection, creating class-specific experts is non-trivial, as images often contain multiple objects from different categories, making it difficult to isolate a single class. To tackle this, we adopt a stochastic training strategy for specializing the $A$ matrices: for each image, we randomly select one of the present classes and update its corresponding $A$ matrix. However, preliminary experiments show this approach is ineffective for rare categories, as random selection limits their exposure to sufficient training data.
@@
Warmup. At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig. 2, yellow arrow) to capture class-agnostic yet task-specific knowledge. This phase, run for a fixed number of epochs, provides a common initialization for the $A$ modules before specialization. By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [32, 38]. Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Sec. 4.3. Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
Specialization. After the warmup phase, we proceed to the core of DitHub: the specialization phase. To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig. 2), each dedicated to a single class. These branches undergo independent optimization, represented by a blue arrow in Fig. 2, to capture class-specific capabilities. In doing so, for each training image, we employ the stochastic strategy introduced in Sec. 4.1 to randomly select a single expert for training.
Notably, the modular design of our approach effectively handles potential class overlaps across consecutive tasks. In real-world scenarios, classes may reappear in new domains with altered distributions or semantics. For instance, consider the class “dog”, originally learned from real-world camera images, later appearing in a task involving thermal imagery. The goal is to retain prior knowledge and enable a “dog” expert to recognize the entity across domains, despite semantic variation. In this case, the modularity of DitHub enables the selective adaptation of the “dog” module, while minimizing interference with other classes, such as the “axolotl” module.
@@
Input: $T$ tasks each containing $|\mathcal{C}_{t}|$ classes; library $\mathcal{A}\leftarrow\{\}$ of class experts; global $B$.
@@
              $A_{c}^{\text{cur}}\leftarrow$ use Eq. 2 $\triangleright$  Merge
@@
     $B\leftarrow$ use Eq. 3
Figure 3: Memory requirements as the number of classes grows.
To favor selective updates, DitHub performs a fetch operation (Fig. 2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Fig. 2) . For instance, considering the class “dog”, the previously stored expert module $A_{\text{dog}}^{\text{old}}$ is fetched and merged with the current task’s warmup matrix $A_{\text{wu}}$ following a weighted strategy:
Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase. Details on the impact of $\lambda_{A}$ are in Sec. 5.2. Crucially, our merging process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information, resulting in $A_{\text{dog}}^{\text{cur}}$, which serves as the starting point for the specialization phase.
@@ 4.2 Bridging Modularity with Memory Efficiency
To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class. As shown in Fig. 3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows. Sharing the $B$ matrix reduces this requirement by half  [^3] In Sec. A.2, we present DitHub’s results when $B$ is not shared.. Additionally, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder alone (line [Uncaptioned image]). This makes DitHub’s memory demands manageable for practical applications. Moreover, memory efficiency is further optimized by fine-tuning a targeted portion of the base architecture (Sec. A.1).
Due to the shared design, the matrix $B$ is more susceptible to interference and catastrophic forgetting [30]. Consequently, we implement a mechanism that ensures smooth updates as tasks progress. Specifically, at the end of each task $t$, we update the matrix $B$ by merging the previous checkpoint, $B_{t-1}$, with the weights learned after the current task, $B^{\text{opt}}$. This process is formulated as follows:
@@
Unlike Eq. 2, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state. This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Sec. 5.2.
4.3 Inference with DitHub
With our library of efficient, class-specific modules in place for adapting the Vision-Language detector, DitHub enables two complementary inference strategies. For requests targeting individual classes where fine-grained adaptation is needed, the corresponding expert module can be directly selected and used to detect that class. Conversely, when evaluating on a textual prompt that includes multiple classes, we retrieve the corresponding class-specific modules from the library and merge them by computing an average of their weights to create a single composite adaptation module. This unified module is then used in a single forward pass to perform detection over all target classes. In our experimental studies, we present results and discussions that address both scenarios.